# The global communication pathways of the human brain transcend the cortical - subcortical - cerebellar division

### v1

@@ Abstract
Neural communication across the cortex, subcortex, and cerebellum is orchestrated by the structural connectome, forming the indispensable anatomical framework for capabilities spanning from elementary motor actions to higher cognitive functions. Yet, despite this importance, the core organizational rules that govern this connectivity remain insufficiently understood. Here we show, for the first time, how the integrated cortical, subcortical, and cerebellar brain areas shape the structural architecture of the whole brain. We find dense structural clusters, which differ in composition and arrangement, vertically transverse the canonical cortical, subcortical, and cerebellar boundaries. These clusters are centralized by a global rich club of predominantly subcortical, alongside cortical hub regions. Congruently, we find that subcortical hubs are not only the most widely connected brain areas but are also leading overall structural integration. Nearly all larger subcortical structures encompass these hub regions, but they also exhibit brain regions with fewer but more specialized connections, pointing toward functional heterogeneity in these structures themselves. Our findings move beyond traditional cortico-centric analysis, offering an initial and global perspective for understanding overall structural connectivity.
@@ Introduction
Complex cognition emerges from the coordinated activity of distributed brain regions. Yet, a historically dominant corticocentric perspective largely confined higher functions to the neocortex, and assigned supportive roles to the subcortical and cerebellar structures Parvizi_Corticocentric_2009 ; Pessoa_Understanding_2014 . Converging evidence now critically challenges this sentiment, demonstrating for example, crucial contributions of the basal ganglia to cognition Obeso_BasalGanglia_2014 , of the hippocampus to decision-making and reward Attaallah_Hippocampus_2024 , and of the cerebellum to cognitive operations Palesi_Pathways_2015 ; Palesi_Contralateral_2017 . To embrace this emerging holistic view of brain function requires a shift towards understanding brain-wide integration, and thus, to consider its communication pathways at the system’s level.
@@
Studies of the white-matter connectivity at the large-scale in the mammalian brain (often referred as the structural connectivity or connectome) have revealed that this forms a hierarchically modular network, centralized by a set of hub regions (termed as the ‘rich-club’) Zamora_Hubs_2010 ; Zamora_FrontReview_2011 ; Heuvel_HubsHuman_2011 . The functional relevance of these hubs has been empirically corroborated for multisensory and cognitive integration Senden_RichClub_2014 ; Senden_RichClub_2017 ; Bertolero_Modular_2015 , and their dysfunction associated with multiple conditions Crossley_Hubs_2014 . So far, structural studies have mostly focused on cortico-cortical connections, ignoring the cerebellum. When included, subcortical structures are often treated as a single node, introducing an unbalanced combination of resolutions. For example, the thalamus is known to broadly connect with the cortex but these projections originate from distinct thalamic nuclei Kumar_Thalamic_2023 . Therefore, considering the thalamus as a single node overestimates its network relevance and it affects computational analyses. On the other hand, detailed examinations of cortico-subcortical links have been reported using either tractography or functional connectivity, but usually following a seed-based approach and focusing on one structure or area, e.g, for thalamus Kumar_Thalamic_2023 ; Bell_Subcortical_2016 , basal ganglia Bell_Subcortical_2016 or brainstem Hansen_Brainstem_2024 . Motivated by these studies, the central role of some subcortical structures to the brain?s communication has been proposed Bell_Subcortical_2016 but, so far, no exhaustive investigation has been undertaken to describe the architecture of the brain-wide connectome, nor to detail how cortex, subcortical structures and cerebellum are interconnected at the macroscopic level in the human brain.
@@
Here, we investigated the human structural connectome including the cortex, the cerebellum and nine subcortical structures at the level of interconnected regions of interest (ROIs), as defined from diffusion imaging and tractography VanEssen_Development_2018 ; Ji_Mapping_2019 . With the entire brain parcellated into 718 ROIs, this network presented sufficient resolution to explore the contribution of individual ROIs even within subcortical structures. Our analyses revealed that the brain-wide connectome follows a modular and hierarchical architecture mirroring the one reported for cortico-cortical connectivity Scannell1993 ; Scannell1995 ; Hilgetag_Clusters_2000 ; Hilgetag_Clustered_2004 ; Zamora_Hubs_2010 ; Zamora_FrontReview_2011 ; Heuvel_HubsHuman_2011 ; Betzel_Modular_2017 ; Suarez_Taxonomy_2022 ; Puxeddu_Relation_2024 , but with significant differences. The human brain is organised into macroscopic network modules, made of intermixed cortical, subcortical and/or cerebellar ROIs thus transcending their individual boundaries. The pathways between these modules are centralised through a communication core (a rich-club) made of cortical and subcortical regions but, importantly, this global rich-club is dominated by subcortical ROIs instead of cortical areas. Surprisingly also, in contrast to previous expectations placing selected structures at the center of the network (e.g., thalamus or basal ganglia), we found all subcortical structures to take part in the rich-club. This is possible because, internally, the ROIs of subcortical structures exhibit heterogeneous network functionalities: some of their ROIs are locally connected while few others are integrated in the global rich-club. Altogether, our study extends the current understanding of the human brain’s organization. It introduces a novel and complementary view by showing how, shaped by white-matter pathways, the brain forms an intermixed network architecture on top of the established division into cortex, subcortex and cerebellum
@@ Results
Structural connectivity data was obtained via diffusion-weighted imaging for 32 healthy human participants from the Human Connectome Project (HCP), see Ref. [VanEssen_Development_2018, ], and parcellated into 718 regions of interest (ROIs). Of them, 360 correspond to cortical, 233 to subcortical, and 125 to cerebellar cortex (see Ref. [Ji_Mapping_2019, ] for parcellation details). This parcellation provides the internal division of subcortical structures into several ROIs such as bilaterally the nucleus accumbens (n=13), brainstem (n=47), caudate nuclei (n=17), diencephalon (n=40), hippocampus (n=29), globus pallidus (n=20), putamen (n=18), thalamus (n=38), and amygdala (n=11). Tractography data was filtered in order to exclude tracts for which the lack of true anatomical fibers has been reported Ramnani_CorticoCereb_2006 ; Palesi_Pathways_2015 ; Palesi_Contralateral_2017 ; DAngelo_Physiology_2018 . Accordingly, the following false positives were filtered out from the connectivity matrix: (i) direct cortico-cerebellar tracks, (ii) direct inter-hemispheric cerebellar connections, and (iii) ipsilateral tracts from the cerebellum to both the thalamus and the brainstem. A population averaged and binary structural adjacency matrix was estimated by applying a distance-aware consensus algorithm. For the rest of the paper, we investigate the properties of this population-level connectivity.
@@
Figure 1: Modular organization of the brain-wide connectome. Structural connectivity matrix and corresponding intra-/inter-modular connection densities for the brain-wide connectivity sorted according a, to the traditional anatomical division of the brain into cortex, subcortex and cerebellum, and b, to the optimal network division into five communities found by community detection algorithms. The optimal network division shows larger intramodular densities and lower cross-modular connection probabilities than for the anatomical division. c, Topographical distribution of the five network communities shows that communities 1-4 vertically stretch over cortical, subcortical, and cerebellar boundaries; communities 3 and 4 are lateralized comprising cortical and subcortical regions and the 5th is made of medially arranged cortical ROIs. d, Quantitative composition of the five network communities in terms of cortical, subcortical or cerebellar ROIs, with hemispheric information added to highlight the lateralization of communities 3 and 4.
We begin by examining the extent to which the network organization of the structural connectivity reflects the traditional anatomical division of the brain into cortex, subcortex and cerebellum. The density matrix in Fig. 1a shows the average probability of an ROI to connect to other ROIs within (diagonal entries) or outside (extradiagonal entries) their anatomical component. Internally, the subcortex is the densest (32%) followed by the cortex (23%) and the cerebellum (23%). In general, we see that brain regions are 1.82 times more likely connected internally, i.e., to other ROIs in their anatomical component than outside. While this may indicate that the anatomical view into cortex, subcortex and cerebellum is a reasonable division from a network point of view, the lack of direct fibers between cortex and cerebellum magnifies this ratio. In fact, cortical ROIs are equally likely connected with each other as to subcortical ROIs (cor-cor: 0.23; cor-sub: 0.24); and cerebellar ROIs are also highly likely connected to subcortical ones (cer-cer: 0.23; cer-sub: 0.19). This observation begs the question whether this anatomical division is also to be expected as the most representative description from the network perspective, or instead, another system?s level organization may be hidden in the connectivity.
@@
We apply a network community detection algorithm (Leidenalg Traag_Leidenalg_2019 ) to automatically cluster the 718 ROIs based on network properties alone. These methods aim at grouping together nodes that are more densely connected with each other than with the nodes in other groups, by maximizing an intra-/inter-connectivity cost function (e.g., Newman modularity). The optimal solution (resolution parameter $\gamma=1.0$ , $Q=0.232$; see Supplementary Fig.  S1) returned the global network divided into five communities (modules or clusters). The adjacency matrix is shown in Fig. 1b reorganized for this optimal partition, alongside its corresponding density matrix. This network-based division clusters the brain more accurately than the anatomical division. The diagonal entries now display notably larger internal connection probabilities (0.34 – 0.51) and lower cross-modular connection probabilities (0.08 – 0.19). On average, under this network-based division, ROIs are 3.23 times more likely connected internally than to other communities.
@@
The five network communities are made of a mixture of regions surpassing the cortical, subcortical, and cerebellar boundaries (Figs. 1c,d; Supplementary Table 1). Community C1 is made of left and right, ventrally located subcortical (43%) and cerebellar ROIs (57%). As seen, the entire cerebellum is embedded into C1 while the Cortex is not present in C1. Community C2 also comprises cortical (44%) and subcortical (56%) ROIs on both hemispheres. Notably, the cortical regions participating in C2 are located in the frontal and posterior lobes, Fig. 1c. Communities C3 and C4 are predominantly lateralized (right and left, respectively), made of cortical and deeper subcortical ROIs. Finally, community C5 (which is the smallest one) is medially located comprising mainly left and right cortical ROIs (96%) and a few subcortical ROIs (4%). These results evidence that communities C1 to C4 are made of regions that transcend the cortical, subcortical, and cerebellar boundaries; bringing up the question of which role do the various subcortical structures play in the cohesion of these macroscopic, intermixed network communities.
@@
Figure 2: Relation between anatomical and optimal network communities. a, Composition of the five optimal network communities in terms of anatomical components, including the division of the subcortex into nine nuclei. Bars indicate the number of ROIs from each anatomical component that was classified into the corresponding network community. Numbers at the bottom summarise the size of the network communities, in total number of ROIs. b, Internal division of the nine subcorticalstructures, the cortex and the cerebellum in terms of the five network-based communities. Topographical maps show that—in most cases except for Accumbens and cerebellum—the ROIs of an anatomical component usually fall into different communities, evidencing an internal distribution of anatomical structures and their participation into different circuits. Bars summarize the number of ROIs that each anatomical component devotes to distinct network communities (C1 – C5). Gray numbers indicate the size (in number of ROIs) of the anatomical components.
The detailed composition of the five network communities is presented in Fig. 2a. C1 is dominated by the cerebellar cortex, the Brainstem and the Dienchephalon, followed by contributions from Pallidum and Hippocampus. Community C2 is made of frontal and posterior cortical ROIs, and a mixture of subcortical nuclei: mainly Thalamus, Hippocampus, Accumbens, Caudate and Amygdala, and to a lesser extent, also the Putamen, Dienchephalon and Pallidum. Communities C3 and C4 are primarily made of lateralized cortical regions together with some subcortical ROIs from Brainstem, Thalamus, Putamen, Pallidum and Caudate. Finally, C5 is practically a cortical community.
Given that the subcortical structures are smaller than the cortex and the cerebellum in number of ROIs, a relative interpretation of these results shall be considered. Hence, we evaluated the fraction of ROIs that each anatomical component contributes to C1 – C5, Fig. 2b. The Putamen is the most segregated subcortical structure, with its 18 ROIs very much distributed among C1 – C4. On the opposite side, the accumbens and the amygdala, both contribute (almost) exclusively to a single community: C2. The rest of nuclei fall in an intermediate situation, with most of their ROIs assigned to one community but few ROIs to other communities.
The involvement of subcortical nuclei into distinct network communities is relevant because it implies that despite their small number of ROIs subcortical structures are internally diverse; with different ROIs potentially participating in different functional circuits. So far, we have analysed the overall division of the brain into network communities and their composition.
@@
Figure 3: Contribution of anatomical components to cross-modular connectivity. a, b, Visualization of the structural connectivity at the level of anatomical components comprising the cortex, the cerebellum and nine subcortical structures; a for the brain-wide network and b the subnetworks representing network communities C1 to C4. Link widths reflect the number of ROIs connected within the two components. c, Illustration of different roles that nodes can play within a modular and hierarchical network, and their characterization. Nodes can be either locally connected inside a module (yellow dot), be well rooted in a community but share connections to other communities (purple triangle), or uniformly participate over all communities (dark purple diamond). d, Mapping of the roles of all the ROIs, considering the network is divided into cortex, subcortex and cerebellum. For visual clarity, the 718 ROIs are presented in three separate maps for cortical, subcortical and cerebellar regions. The solid horizontal lines represent the average node degree and gray shadows the range for one standard deviation. e, Mapping of the roles for all the ROIs, considering the network is divided into the five network communities. Top left panel shows all the 718 ROIs on a single map. The remaining panels show the same results but individually for the ROIs of each of the 11 anatomical components studied.
@@
Next, we investigate how the identified brain-wide communities are interconnected and explore, specifically, the contribution of subcortical connectivity. Network visualization algorithms allow for an approximate representation of a network’s architecture by arranging their nodes according to their relative importance and mutual similarity. The whole-brain connectome is displayed in Fig. 3a (Force Atlas 2 algorithm Jacomy_ForceAtlas_2014 . In this visualization the network appears centered around the two cortical hemispheres (in green) largely connected to the Thalamus and the Putamen on the one hand, and the Brainstem, the Diencephalon and the Pallidum on the other hand mediating the cortico-cerebellar pathways. The smaller subcortical structures tend to be placed around the periphery, with those contributing mostly to the C2 community (e.g., Accumbens, Amygdala, Hippocampus and the Thalamus; see Fig. 3b) occupying the upper space while the structures more associated to C1 (e.g., Brainstem, Diencephalon and Pallidum; Fig. 3b) are placed towards the bottom, near the Cerebellum.
@@
This spatial separation is reflected in more detail for the subnetworks formed by the individual network communities for C1 and C2, Fig. 3b. The core of C1 is made of substantial interconnectivity between the Cerebellum, the Brainstem (specially its central part) and the Diencephalon. Instead, C2 is governed by strong interconnections between the two cortical hemispheres and a substantial number of connections to the Hippocampus, the Thalamus and the Caudate. C3 and C4 form star-like structures centered on either right or left cortical ROIs.
@@
Although these layouts are informative, they are affected by the size differences between brain regions. Looking at connection probabilities instead of total link numbers, some smaller areas like the Putamen, Amygdala, and Thalamus become more central, highlighting their key role in connecting different parts of the brain (see Supplementary Figure S2).
@@
We now clarify in more detail the position that every ROIs takes within the brain-wide connectome. For that, we map the 718 ROIs based on two parameters Guimera_RolesJSM_2005 ; Zamora_FrontReview_2011 ; Klimm_Roles_2014 : their degree (i.e., the hubness) and their participation index (which quantifies how evenly the connections of an ROI are distributed across the modules). See Fig. 3c for a schematic description. Hence, the participation of a node depends on how the network was divided into modules in the first place.
@@
Considering the anatomical division (Fig. 3d), cortical and cerebellar ROIs tend to occupy the lower-right and the central sectors of the map, corresponding to nodes with local or intermediate cross-modular connections. The reason is the absence of direct connections between Cerebellum and Cortex. Thus, under this anatomical view, cortical and cerebellar ROIs span their connections, at most, among two of the three modules (i.e., either cortex and subcortex, or cerebellum and subcortex) and thus attain a maximal participation index of 0.5. On the other hand, several subcortical ROIs occupy the upper-right sector of the map which is characteristic of supramodal hubs (e.g., purple diamond in Fig. 3c). Importantly, this indicates that at the systems? level, the brain-wide communication is primarily sustained by subcortical regions.
@@
Considering the network-based division (Fig. 3e), the 718 ROIs distribute along the diagonal axis of the 2D map (top-left panel), which is a signature of a modular and hierarchical organization Zamora_FrontReview_2011 ; Klimm_Roles_2014 ; Zamora_FComplexity_2016 . Disentangling the results individually by anatomical component, the ROIs of the cortex display a wider variety of roles as compared to the results under the anatomical division. It now includes some hubs of high participation (up to 0.86). On the contrary, cerebellar ROIs still fall into the lower-left sectors of the map. Given that C3 – C5 are mainly made of cortical regions and the absence of direct cortico-cerebellar fibers, cerebellar ROIs span their connections mainly among C1 and C2. For the subcortical structures, the results show a rich variety of connectivity roles. For example, all ROIs of the Accumbens (n=13) and the Amygdala (n=11) were classified into C2, but some of their ROIs take a central role with large degree and high participation. The Caudate (n=17) and the Putamen (n=18) are two other small nuclei that contain ROIs with high degree and participation, although in this case it is more expected, given that they contribute to C2 – C5 and C1 – C4 respectively, Fig. 2.
@@
Altogether, these results show that subcortical structures are internally heterogeneous. Within each structure, their ROIs occupy different positions in the network architecture thus playing differential roles for the cross-modular communication. Some ROIs form mostly local connections (characterised by low degree and participation) and others span their connections across multiple communities (large degree and participation) becoming crucial for the centralization of the brain-wide pathways.

### v2

@@ Abstract
Understanding how cortex, subcortex and cerebellum integrate is a major challenge for neuroscience, however, studies of the brain’s structural connectivity have mostly focused on cortico-cortical links. Here, we used diffusion imaging to construct the structural connectome of the entire human brain including 360 cortical, 233 subcortical, and 125 cerebellar regions of interest (ROIs). We found that the brain forms a modular and hierarchical network architecture, organized into modules of mixed cortical, subcortical and/or cerebellar regions, and whose cross-modular pathways are centralized through highly connected hub ROIs (a ‘rich-club’). This global rich-club is subcortically dominated and, surprisingly, composed of hub ROIs from all subcortical structures rather than one region like the thalamus, centralizing the communication pathways. This study improves our understanding of the human brain’s organization. It provides structural evidence to question the prevalent cortico-centric notion by revealing a connectome centered at the subcortex but made of transversal pathways.
@@ Introduction
Complex cognition emerges from the coordinated activity of distributed brain regions. Historically, a dominant cortico-centric perspective of the brain largely confined higher functions to the neocortex, while subcortical and cerebellar structures were considered to play supportive roles Parvizi (2009); Pessoa (2014). Converging evidence now critically challenges this conception by demonstrating, for example, the crucial contribution of the basal ganglia to cognition Obeso et al. (2014), of the hippocampus to decision-making and reward Attaallah et al. (2024), and of the cerebellum to cognitive operations Palesi et al. (2015, 2017). An important missing gap to fully embrace this emerging holistic view of brain function is to understand how the cortex, subcortex and cerebellum communicate with each other at the systems level.
@@
The study of structural connectivity derived from white-matter fibers has revealed that the brain is shaped into a hierarchically modular network, centralized by a set of hub regions termed the ‘rich-club’ Zamora-López et al. (2010, 2011); van den Heuvel and Sporns (2011). The functional relevance of these hubs has been empirically corroborated for multi-sensory and cognitive integration Senden et al. (2014, 2017); Bertolero et al. (2015), and their dysfunction has been associated with various neuropsychiatric diseases Crossley et al. (2014). So far, these connectivity studies have excluded the cerebellum and, when included, the subcortical structures were treated as a single node of the network introducing an unbalanced mixture of resolutions. For example, the thalamus broadly connects with the cortex but these projections originate from distinct subthalamic nuclei Jones (2001); Kumar et al. (2023). Therefore, considering the whole thalamus as a single node overestimates its network relevance and affects the outcome of computational models. On a local scale, detailed descriptions of cortico-subcortical links have been reported using either tractography or functional connectivity, but usually focusing on the connectivity of one structure alone, such as the thalamus Kumar et al. (2023); Bell and Shine (2016), basal ganglia Bell and Shine (2016) or brainstem Hansen et al. (2024). Accordingly, while the central role of subcortical structures has been highlighted, no exhaustive investigation has yet been undertaken to describe how the various subcortical structures are interconnected at the macroscopic level nor to clarify the place each takes in the integrated architecture of the entire brain.
@@
Here, we employed diffusion imaging and tractography to investigate the structural architecture of the human brain encompassing the cortex, the cerebellum and nine subcortical structures. With the brain parcellated into 718 regions of interest (ROIs), this network presented sufficient resolution to separate subcortical structures into multiple ROIs allowing to explore their detailed connectivity. Our analyses revealed that this brain-wide human connectome follows a modular and hierarchical architecture mirroring the one reported for cortico-cortical connectivity alone Scannell and Young (1993); Scannell et al. (1995); Hilgetag et al. (2000); Hilgetag and Kaiser (2004); Zamora-López et al. (2010, 2011); van den Heuvel and Sporns (2011); Betzel et al. (2017); Suarez et al. (2022); Puxeddu et al. (2024), but with significant differences. First, the human brain is organized into macroscopic network modules, composed of intermixed cortical, subcortical and/or cerebellar ROIs thus transcending their individual boundaries. Second, the pathways between these modules are centralized through a communication core (a ‘rich-club’) that contains cortical and subcortical regions but, importantly, is dominated by subcortical rather than by cortical ROIs. In fact, we showed that in-silico lesions of subcortical hubs greatly disrupted the communication throughout the network, significantly more than the lesion of cortical hubs. Last, and surprisingly, we found subcortical structures to be internally heterogeneous. Their ROIs show a diversity of local and global connectivity profiles indicating they play differential functional roles. A relevant consequence of this internal heterogeneity is that all subcortical structures contribute—partly—to the global rich-club, in contrast to previous reports prone to place one structure (e.g., thalamus or basal ganglia) at the center of the brain-wide communication. Altogether, our study improves the current understanding of the human brain’s organization. It introduces a novel and complementary view, providing structural arguments to dispute the prevalent cortico-centric notion that regards subcortical structures and cerebellum as mere supporters of cortical function. Instead, our results reveal a network architecture centered at the subcortex but made of transversal pathways.
@@ Results
Structural connectivity data was obtained via diffusion-weighted imaging for 32 healthy human participants from the Human Connectome Project (HCP) van Essen et al. (2018) and parcellated into 718 regions of interest (ROIs). Of these, 360 correspond to cortical, 233 to subcortical, and 125 to cerebellar cortex (see Ji et al. (2019) for details). This parcellation provides the internal division of subcortical structures into several ROIs such as the bilateral nucleus accumbens (n=13), brainstem (n=47), caudate nuclei (n=17), diencephalon (n=40), hippocampus (n=29), globus pallidus (n=20), putamen (n=18), thalamus (n=38), and amygdala (n=11). Tractography data was filtered in order to exclude false-positive tracts related to cerebellar connections Ramnani (2006); Palesi et al. (2015, 2017); D’Angelo (2018). A population averaged, binary structural adjacency matrix was estimated employing a distance-aware consensus algorithm. For the rest of the paper, we investigate the properties of this population-level connectivity.
@@
Figure 1: Modular organization of the brain-wide connectome. Structural connectivity matrix and corresponding intra-/inter-modular connection densities for the brain-wide connectivity sorted according a, to the traditional anatomical division of the brain into cortex, subcortex and cerebellum, and b, to the optimal network division into five communities found by community detection algorithms. The optimal network division shows larger intra-modular densities and lower cross-modular connection probabilities than for the anatomical division. c, Topographical distribution of the five network communities shows that communities 1-4 vertically stretch over cortical, subcortical, and cerebellar boundaries; communities 3 and 4 are lateralized comprising cortical and subcortical regions and the 5th is made of medially arranged cortical ROIs. d, Quantitative composition of the five network communities in terms of cortical, subcortical or cerebellar ROIs, with hemispheric information added to highlight the lateralization of communities 3 and 4.
We begin by inquiring whether the anatomical division of the brain into cortex, subcortex and cerebellum is also reflected in the organization of its structural connectivity. For that, we estimated the average probability that an ROI connects to other ROIs in the same anatomical category (diagonal entries of density matrix in Fig. 1a) and to ROIs of other categories (extradiagonal entries). In terms of network architecture, a good division requires intra-category probabilities to be notably larger than inter-category probabilities. However, we see that cortical ROIs are equally likely connected with each other as with subcortical ROIs (cor-cor: 0.23; cor-sub: 0.24), while cerebellar ROIs are very likely connected to subcortical ones too (cer-cer: 0.23; cer-sub: 0.19). This observation indicates that this traditional, anatomical classification does not represent a faithful organization of the brain’s connectivity into network modules.
@@
To identify a more appropriate description of the brain’s systems level organization, we apply a community detection algorithm (Leidenalg Traag et al. (2019)) to automatically cluster the 718 ROIs based on network properties alone. This algorithm aims at grouping together nodes that are more densely connected with each other than with the nodes in other groups, by maximizing an intra-/inter-connectivity cost function (e.g., Newman modularity). The optimal solution (resolution parameter $\gamma=1.0$, $Q=0.232$; see Supplementary Fig. S1) returned a division of the brain-wide connectome into five communities (modules or clusters). The adjacency matrix is shown in Fig. 1b, reorganized for this optimal partition alongside its corresponding density matrix. This network-based division clusters the brain notably better than the anatomical division (Fig. 1a). The diagonal entries now display larger internal connection probabilities (0.34 – 0.51) and lower cross-modular connection probabilities (0.08 – 0.19). On average, under this network-based division, ROIs are 3.23 times more likely connected to other ROIs in their community than with ROIs in other communities.
@@
The five network communities consist of a mixture of regions surpassing the cortical, subcortical, and cerebellar boundaries (Figs. 1c,d). Community C1 contains left and right, ventrally located subcortical (43%) and cerebellar ROIs (57%). The entire cerebellum is embedded into C1. Community C2 comprises cortical (44%) and subcortical (56%) ROIs on both hemispheres. Notably, the cortical regions participating in C2 are located in frontal and posterior lobes, Fig. 1c. Communities C3 and C4 display strong lateralization (right and left) and are predominantly composed of cortical ROIs (84% and 86%), with fewer subcortical ROIs (16% and 14%). Finally, a smaller community C5 is medially located comprising left and right cortical ROIs (96%) and a few subcortical ROIs (4%). These results provide evidence that, from a network perspective, the brain-wide connectome is organized into network modules that transcend the cortical, subcortical, and cerebellar boundaries.
@@
Figure 2: Relation between anatomical and optimal network communities. a, Composition of the five optimal network communities in terms of anatomical components, including the division of the subcortex into nine nuclei. Bars indicate the number of ROIs from each anatomical component that was classified into the corresponding network community. Numbers at the bottom summarize the size of the network communities, in total number of ROIs. b, Internal division of the nine subcortical structures, the cortex and the cerebellum in terms of the five network-based communities. Topographical maps show that—in most cases except for Accumbens and cerebellum—the ROIs of an anatomical component usually fall into different communities, evidencing an internal distribution of anatomical structures and their participation into different circuits. Bars summarize the number of ROIs that each anatomical component devotes to distinct network communities (C1 – C5). Gray numbers indicate the size (in number of ROIs) of the anatomical components.
We now take a closer look at the composition of these five network communities to better identify the role of subcortical structures in this division, Fig. 2a. C1 is mainly made of ROIs from the cerebellum, brainstem and diencephalon, followed by contributions from pallidum and hippocampus. Community C2 is made of frontal and posterior cortical ROIs, and a mixture of subcortical nuclei: mainly thalamus, hippocampus, accumbens, caudate and amygdala, and to a lesser extent, also the putamen, diencephalon and pallidum. Communities C3 and C4 are made of lateralized cortical regions together with subcortical ROIs from brainstem, thalamus, putamen, pallidum and caudate. Last, C5 is practically a cortical community.
Given that the subcortical structures are smaller than cortex and cerebellum (in number of ROIs), a relative interpretation of these results is needed. Thus, Fig. 2b shows how the ROIs of each structure are devoted to the different communities. The putamen is the most segregated subcortical structure with its 18 ROIs very much distributed among C1 – C4. On the opposite side, the accumbens and the amygdala contribute (almost) all their ROIs to C2. The remaining structures show an intermediate situation, with most of their ROIs assigned to one community and fewer ROIs contributing to other communities.
The involvement of subcortical structures in distinct network communities is relevant because it implies that despite their small size, subcortical structures are internally diverse, and indicates their participation in many functional circuits.
@@
Figure 3: Contribution of anatomical components to cross-modular connectivity. a, b, Visualization of the structural connectivity at the level of anatomical components comprising the cortex, the cerebellum and nine subcortical structures; a, for the brain-wide network and b, the subnetworks representing network communities C1 to C4. Community C5 is omitted for its simplicity as it only contains three regions. Node diameters are proportional to the number of ROIs to each anatomical component. Link widths reflect the number of ROIs connected within the two components. c, Illustration of different roles that nodes can play within a modular and hierarchical network, and their characterization. Nodes can be either locally connected inside a module (yellow dot), be well rooted in a community but share connections to other communities (purple triangle), or uniformly participate over all communities (dark purple diamond). These differential contributions are typically characterized by two parameters allowing to project each node into a 2D map (axes on the right panel). (i) The degree, informing of the overall centrality of a node and (ii) the participation index, indicating how a node spreads its links among communities. Participation 0.0 means the node is only locally connected in one module (e.g., the blue dot) while participation 1.0 means a node is equally likely connected to all communities (e.g., the purple diamond). Locally connected nodes are expected to play more specialized functions while hubs with large participation are usually associated with multifunctional roles due to their access to a wider range of information. d, Mapping of the roles of all the ROIs, considering the network is divided into cortex, subcortex and cerebellum. For visual clarity, the 718 ROIs are presented in three separate maps for cortical, subcortical and cerebellar regions. The solid horizontal lines represent the average node degree and gray shadows the range for one standard deviation. e, Mapping of the roles for all the ROIs, considering the network is divided into the five network communities. Top left panel shows all the 718 ROIs on a single map. The remaining panels show the same results but individually for the ROIs of each of the 11 anatomical components studied. ROIs are colored according to the network community they were assigned to.
@@
So far, we have explored the overall division of the brain into network communities and their composition. Now, we investigate how these brain-wide communities connect with each other. Network visualization algorithms allow for an approximate representation of a network’s architecture by arranging their nodes according to their relative importance and mutual similarity. The whole-brain connectome is displayed in Fig. 3a (Force Atlas 2 algorithm Jacomy et al. (2014)). In this visualization the network appears centered around the two cortical hemispheres (in green) largely connected to the thalamus and the putamen on the one hand, with the brainstem, diencephalon and pallidum mediating the cortico-cerebellar pathways on the other hand. The smaller subcortical structures tend to be placed around the periphery, with those contributing mostly to the C2 community (e.g., accumbens, amygdala, hippocampus and thalamus; see Fig. 2b) occupying the upper space while the structures more associated to C1 (e.g., brainstem, diencephalon and pallidum; Fig. 2b) are placed towards the bottom, near the cerebellum.
@@
This spatial separation is reflected in more detail when observing the subnetworks formed individually by the network communities C1 and C2, Fig. 3b. The core of C1 exhibits substantial interconnectivity between the cerebellum, the brainstem (specially its central part) and the diencephalon. Instead, C2 is governed by strong interconnections between the two cortical hemispheres and a substantial number of connections to the hippocampus, thalamus and caudate. C3 and C4 form star-like structures centered on either right or left cortical ROIs.
@@
Although these layouts are informative, they are affected by the size differences (in number of ROIs) of the anatomical components. Looking at connection probabilities instead of total link numbers, some smaller areas like the putamen, amygdala, and thalamus become more central, highlighting their key role in connecting different parts of the brain (see Supplementary Fig. S2).
@@
We now clarify in more detail the position that every ROIs takes within the brain-wide connectome. For that, we map the 718 ROIs based on two parameters Guimerà and Amaral (2005); Zamora-López et al. (2011); Klimm et al. (2014): their degree (i.e., hubness) and their participation index which quantifies how (un-)evenly the connections of an ROI are distributed across modules. See Fig. 3c for a schematic description. Hence, the participation index of a node depends on how the network was divided into modules in the first place.
@@
Considering the anatomical division (Fig. 3d), cortical and cerebellar ROIs tend to occupy the lower-right and the central sectors of the map, corresponding to nodes with local or intermediate cross-modular connections. The reason is the absence of direct connections between cerebellum and cortex. Thus, under this anatomical view, cortical and cerebellar ROIs span their connections, at most, among two of the three modules (i.e., either cortex and subcortex, or cerebellum and subcortex) and thus attain a maximal participation index of 0.5. On the other hand, several subcortical ROIs occupy the upper-right sector of the map which is characteristic of supramodal hubs (e.g., purple diamond in Fig. 3c). Importantly, this also indicates that the brain-wide communication is mainly sustained by subcortical regions.
@@
Considering the network-based division (Fig. 3e), the 718 ROIs distribute along the diagonal axis of the 2D map (top-left panel), which is a signature of a modular and hierarchical organization Zamora-López et al. (2011); Klimm et al. (2014); Zamora-López et al. (2016). Disentangling the results individually by anatomical structures, the ROIs of the cortex display a wider variety of roles as compared to the results under the anatomical division. It now includes some hubs of high participation (up to 0.86). On the contrary, cerebellar ROIs still fall into the lower-left sectors of the map, as in fact, cerebellar ROIs span their connections mostly among ROIs in C1 and C2. For the subcortical structures, the results show a rich variety of connectivity roles. For example, the accumbens and the amygdala were entirely classified into community C2, however, some of their ROIs display both large degree and high participation, indicating their connections span throughout all communities and thus taking a central role in the network. The caudate and the putamen are two other small nuclei that contain ROIs with high degree and participation, although in this case it was more expected given that their ROIs were distributed among C2 – C5 and C1 – C4 respectively, Fig. 2.
@@
Altogether, these results show that subcortical structures are internally heterogeneous. Their ROIs occupy different positions in the degree-participation maps, thus indicating they play differential roles in the network architecture. Some ROIs form mostly local connections (characterized by low degree and participation) and others span their connections across multiple communities (large degree and participation) becoming crucial for the centralization of the brain-wide pathways.